# Partial Functional Dynamic Backdoor Diffusion-based Causal Model

### v1

@@ Abstract
We introduce a Partial Functional Dynamic Backdoor Diffusion-based Causal Model (PFD-BDCM), specifically designed for causal inference in the presence of unmeasured confounders with spatial heterogeneity and temporal dependency. The proposed PFD-BDCM framework addresses the restrictions of the existing approaches by uniquely integrating models for complex spatio-temporal dynamics with the analysis of multi-resolution variables. Specifically, the framework systematically mitigates confounding bias by integrating valid backdoor adjustment sets into a diffusion-based sampling mechanism. Moreover, it accounts for the intricate dynamics of unmeasured confounders through the deployment of region-specific structural equations and conditional autoregressive processes, and accommodates variables observed at heterogeneous resolutions via basis expansions for functional data. Our theoretical analysis establishes error bounds for counterfactual estimates of PFD-BDCM, formally linking reconstruction accuracy to counterfactual fidelity under monotonicity assumptions of structural equation and invertibility assumptions of encoding function. Empirical evaluations on synthetic datasets and real-world air pollution data demonstrate PFD-BDCM’s superiority over existing methods.
Keywords: Unmeasured Confounders; Spatiotemporal Dynamic Causal Graphs; Diffusion-based Causal Models; Backdoor Criterion
@@ 1 Introduction
In the Structural Causal Model (SCM) framework, causal queries can be answered by learning a proxy for the unobserved exogenous noise and the structural equations (Pearl,, 2009). This suggests that (conditional) generative models that encode to a latent space could be a option for modeling SCMs, as the latent space serves as proxies for exogenous noises. In the SCM, the encoding process extracts the latent space from an observation, and the decoding process generates the sample from the latent space, approximating the structural equations. Chao et al., (2023) generalized their approach to arbitrary causal graphs, and proposed the Diffusion-based Causal Model (DCM) that allowed one to sample from the target distribution of interest, by which it can estimate the average treatment effect (ATE), outperforming some of the state-of-the-art algorithms (Sanchez-Martin et al.,, 2021; Khemakhem et al.,, 2021). DCM assumes causal sufficiency (i.e., absence of hidden confounders) and focused on approximating the structural equations given observational data and the underlying causal directed acyclic graph (DAG), and providing a mechanism for answering the causal observational, interventional and counterfactual queries (Pearl,, 2009). The DCM assumes causal sufficiency, which needs observing all confounding variables, which usually fails in applications. To overcome the limitation, Shimizu, (2023) proposed the BDCM algorithm for estimating ATE in the presence of unmeasured confounders by including nodes that satisfy the backdoor criterion (Pearl et al.,, 2016) in both training and sampling phases.
Both DCM and BDCM are for static causal relationships, whereas real-world causal relationships among variables may be affected by the spatio-temporal unmeasured confounding factors. Therefore, we propose a Partial Functional Dynamic Backdoor Diffusion-based Causal Model (PFD-BDCM), which can uncover the causal relationships among variables across different resolutions in the presence of spatio-temporal unmeasured confounders.
@@
We propose spatio-temporal diffusion models for modeling partial functional spatio-temporal dynamic causal relationships. Diffusion models (Sohl-Dickstein et al.,, 2015; Ho et al.,, 2020; Song et al.,, 2021) have gained prominence due to their high expressiveness and performance in generative tasks (Saharia et al.,, 2022; Ramesh et al.,, 2022; Kong et al.,, 2020). Our primary contribution is to show how to apply diffusion models to capture partially functional dynamic causal relationship with unmeasured confounders. The core idea is in modeling each node in dynamic causal graphs with a spatio-temporal diffusion model and cascading the generated samples in the topological order to answer causal queries. For each node, its corresponding spatio-temporal diffusion model encodes and decodes latent representations using variables of nodes. Within our framework, the encoding and decoding procedures extend the Denoising Diffusion Implicit Models (DDIMs) (Song et al.,, 2021) paradigm by integrating backdoor adjustment sets as supplementary covariates. We term the resulting framework as the Partial Functional Dynamic Backdoor Diffusion-based Causal Model (PFD-BDCM). Theoretical analysis demonstrate that PFD-BDCM preserves essential structural properties of Structural Causal Models (SCMs). Key contributions of our study include:
  1. 1.
@@
    The proposed PFD-BDCM provides a unified framework for approximating both interventions (do-operator) and counterfactuals (abduction-action-prediction steps). It has a training procedure requiring only the dynamic causal graph and observational data, and the trained model enables: i) sampling from observational/interventional distributions; ii) precise counterfactual query resolution.
  2. 2.
@@
    Our theoretical analysis proves that the counterfactual estimates given by PFD-BDCM admit quantifiable error bounds under reasonable assumptions. i) It provides the first formal explanation for the performance gains of encoder-decoder architectures (e.g., diffusion models) in counterfactual querying through error bounds; ii) It extends to the more challenging multivariate case under an additional assumption and to diverse encoder-decoder models.
  3. 3.
    We evaluate the performance of PFD-BDCM in empirical settings involving spatio-temporal dynamic structure equations and three forms of causal queries.
2 Preliminaries
Notations.
@@ 2 Preliminaries
To distinguish between the nodes in the causal graph and diffusion random variables, we use subscripts to denote graph nodes. Let $n]:=\{1,\cdots,n\}$, $\mathrm{dim(x)}$ represents the dimension of $x$.
Structural Causal Models.
@@
Consider a directed acyclic graph (DAG) $\mathcal{G}$ with nodes $[K]$ in a topologically sorted order  [^1] A topological order is a linear arrangement of variables where a variable appears after all its direct causes (parents) (Pearl,, 2009)., where a node $k$ is built with an endogenous random variable $X_{k}$ defined on a space $\mathcal{X}_{k}\subset\mathbb{R}^{d_{k}}$, which has a random exogenous input $U_{k}$. Let $\mathrm{PA}_{k}$ be the set of parent nodes of node $k$ in $\mathcal{G}$ and let $X_{\mathrm{PA}_{k}}:=\left\{X_{l}\right\}_{l\in\mathrm{PA}_{k}}$ be a set of parent random variables defined on $\mathrm{PA}_{k}$.
A structural causal model (SCM) $\mathcal{M}$ characterizes the relationship between the endogenous variable of a node with the endogenous variables of its parents and its own exogenous variable. Formally, we define $\mathcal{M}:=(\mathbf{F}(\mathbf{X},\mathbf{U}),p_{U})$, where $\mathbf{F}(\mathbf{X},\mathbf{U})$ specifies how entire endogenous variables $\mathbf{X}:=\{X_{1},\cdots,X_{K}\}$ are generated from the set of exogenous random variables $\mathbf{U}:=\{U_{1},\cdots,U_{K}\}$ with a prior distribution $p_{U}$. The structural mechanism is governed by $\mathbf{F}(\mathbf{X},\mathbf{U}):=(f(X_{\mathrm{PA}_{1}},U_{1}),\cdots,f(X_{\mathrm{PA}_{K}},U_{K}))$, where each $X_{k}:=f(X_{\mathrm{PA}_{k}},U_{k})$ for $k\in[K]$ (Pearl,, 2009).
Conditional Diffusion Model.
Recent advances in deep learning-based causal inference (Chao et al.,, 2023; Shimizu,, 2023) have demonstrated the effectiveness of diffusion models for answering causal queries. Our approach leverages the generative power of diffusion models to learn the complex functional relationships inherent in a Structural Causal Model. The core idea is to represent the structural equation for each endogenous variable $X_{k}$ with a dedicated conditional diffusion model. This model learns the distribution $p(X_{k}|X_{\mathrm{PA}_{k}})$ .

### v2

# Partially Functional Dynamic Backdoor Diffusion-based Causal Model

@@ Abstract
Causal inference in settings involving complex spatio-temporal dependencies, such as environmental epidemiology, is challenging due to the presence of unmeasured confounding. However, a significant gap persists in existing methods: current diffusion-based causal models rely on restrictive assumptions of causal sufficiency or static confounding. To address this limitation, we introduce the Partially Functional Dynamic Backdoor Diffusion-based Causal Model (PFD-BDCM), a generative framework designed to bridge this gap. Our approach uniquely incorporates valid backdoor adjustments into the diffusion sampling mechanism to mitigate bias from unmeasured confounders. Specifically, it captures their intricate dynamics through region-specific structural equations and conditional autoregressive processes, and accommodates multi-resolution variables via functional data techniques. Furthermore, we provide theoretical guarantees by establishing error bounds for counterfactual estimates. Extensive experiments on synthetic data and a real-world air pollution case study confirm that PFD-BDCM outperforms current state-of-the-art methods.
@@ 1 Introduction
Within the Structural Causal Model (SCM) framework, causal queries can be answered by learning a proxy for the unobserved exogenous noise and the structural equations (Pearl,, 2009). This suggests that (conditional) generative models that encode to a latent space could be an option for modeling SCMs, as the latent space serves as proxies for exogenous noises. Recent advances have explored the integration of deep generative models with structural causal models to address these challenges. Chao et al., (2023) proposed the Diffusion-based Causal Model (DCM), which leverages diffusion processes to approximate structural equations and answer causal queries without explicit intervention data. However, DCM assumes causal sufficiency (no unobserved confounders) which rarely holds in practice. Shimizu, (2023) extended this line of work with Backdoor Diffusion-based Causal Model (BDCM), incorporating backdoor adjustment to handle certain types of unmeasured confounding. Nevertheless, both methods operate under static assumptions and fail to account for spatio-temporal structure in confounding variables, a critical limitation in real-world systems where confounders often exhibit complex dependencies across space and time.
To address this gap, we propose the Partially Functional Dynamic Backdoor Diffusion-based Causal Model (PFD-BDCM), a generative framework capable of explicitly modeling spatiotemporal correlations among unmeasured confounders while supporting causal inference. First, it integrates valid backdoor adjustments into the diffusion sampling mechanism to mitigate bias from unmeasured confounders. Second, it explicitly models the intricate spatio-temporal dynamics of unmeasured confounders through region-specific structural equations and conditional autoregressive processes. Third, it accommodates variables observed at heterogeneous resolutions via functional data techniques and basis expansions. Furthermore, modeling each node to more accurately address causal inference by leveraging the superior data generation capability of diffusion models (Song et al.,, 2021). Our approach relaxes the causal sufficiency assumption and captures dynamic confounding through a structured latent representation, allowing for more robust estimation of causal effects in non-stationary environments. We demonstrate empirically that PFD-BDCM outperforms existing diffusion-based causal models across a range of observational, interventional and counterfactual queries, particularly in settings with spatially and temporally varying confounders.
@@
We propose a Partially Functional Dynamic Backdoor Diffusion-based Causal Model (PFD-BDCM) for modeling partially functional spatio-temporal dynamic causal relationships. Diffusion models (Sohl-Dickstein et al.,, 2015; Ho et al.,, 2020; Song et al.,, 2021) have gained prominence due to their high expressiveness and performance in generative tasks. Our primary contribution is to show how to apply diffusion models to capture partially functional dynamic causal relationship with unmeasured confounders. The core idea is in modeling each node in Partially Functional Spatio-Temporal Dynamic Structural Causal Model (PFST-DSCM) with a diffusion model and cascading the generated samples in the topological order to answer causal queries. For each node, the encoding process of the diffusion model takes the current node variable and its set of backdoor node variables as input to generate latent representations of the unobserved exogenous variables, and then reconstructs the current node variable through a decoding process. Within our framework, the encoding and decoding procedures extend the Denoising Diffusion Implicit Models (DDIMs) (Song et al.,, 2021) paradigm by integrating backdoor adjustment sets as supplementary covariates. Key contributions of our study include:
@@
(1) [Section 2] The proposed PFD-BDCM provides a unified framework for approximating both interventions (do-operator) and counterfactuals (abduction-action-prediction steps). It has a training procedure requiring only the dynamic causal graph and observational data, and the trained model enables: i) sampling from observational/interventional distributions; ii) precise counterfactual query resolution.
@@
(2) [Section 3] Our theoretical analysis proves that the counterfactual estimates given by PFD-BDCM admit quantifiable error bounds under reasonable assumptions. i) It provides the first formal explanation for the performance gains of encoder-decoder architectures (e.g., diffusion models) in counterfactual querying through error bounds; ii) It extends to the more challenging multivariate case under an additional assumption and to diverse encoder-decoder models.
(3) [Section 4] We evaluated the performance of PFD-BDCM on three synthetic datasets of varying scales involving spatiotemporal dynamic structural equations and three types of causal queries. Experimental results demonstrate that PFD-BDCM consistently outperforms existing state-of-the-art methods (Chao et al.,, 2023; Shimizu,, 2023) as well as PFD-DCM. Furthermore, we also demonstrate the strong performance of PFD-BDCM on a real-world atmospheric pollution dataset.
2 Methodology
We first introduce some useful notations and concepts.
@@ 2 Methodology
Notations: To distinguish between the nodes in the causal graph and diffusion random variables, we use subscripts to denote graph nodes. Let $n]:=\{1,\cdots,n\}$ and $\mathrm{dim}(x)$ represents the dimension of $x$; let $Z_{k}^{t}$ be the exogenous noise at diffusion step $t$ in the forward process, with $Z_{k}:=Z_{k}^{T}$, and $\hat{X}_{k}^{t}$ the endogenous variable at step $t$ in the reverse process, where $\hat{X}_{k}:=\hat{X}_{k}^{0}$.
In causal inference, a confounder denotes a variable that causally influences both a treatment variable $X$ and an outcome variable $Y$, thereby inducing a non-causal association between them. (Pearl,, 2009). Observable Confounders refer to confounders that can be measured, which permit adjustment through statistical methods such as stratification, matching or regression (Pearl,, 2009). Unobservable Confounders denote latent variables that fulfill confounding criteria but resist direct measurement, which can potentially bias causal estimates when unaccounted for. Unobserved explanatory (explained) nodes are unobserved confounder nodes that have no unobserved confounder nodes as its parent (descendant).
Figure 1: DAG with three nodes (left) and SCM with three exogenous and endogenous nodes (right)
@@
Structural Causal Models. Consider a directed acyclic graph (DAG) (such as Fig. 1) $\mathcal{G}$ with nodes $[K]$ in a topologically sorted order  [^1] A topological order is a linear arrangement of variables where a variable appears after all its direct causes (parents) (Pearl,, 2009), where a node $k$ is built with an endogenous random variable $X_{k}$ defined on a space $\mathcal{X}_{k}\subset\mathbb{R}^{d_{k}}$, which has a random exogenous input $U_{k}$. Let $\mathrm{PA}_{k}$ be the set of parent nodes of node $k$ in $\mathcal{G}$ and let $X_{\mathrm{PA}_{k}}:=\left\{X_{l}\right\}_{l\in\mathrm{PA}_{k}}$ be the set of parent random variables of $\mathrm{PA}_{k}$. A structural causal model (SCM) $\mathcal{M}$ characterizes the relationship between the endogenous variable $X_{k}$ of a node $K$ with the endogenous variables of its parents $X_{\mathrm{PA}_{k}}$ and its own exogenous variable $U_{k}$. Formally, we define $\mathcal{M}:=(\mathbf{F}(\mathbf{X},\mathbf{U}),p_{\mathbf{U}})$, where $\mathbf{F}(\mathbf{X},\mathbf{U})$ specifies how entire endogenous variables $\mathbf{X}:=\{X_{1},\cdots,X_{K}\}$ are generated from the set of exogenous random variables $\mathbf{U}:=\{U_{1},\cdots,U_{K}\}$ with a prior distribution $p_{\mathbf{U}}$. The structural mechanism is governed by $\mathbf{F}(\mathbf{X},\mathbf{U}):=(f(X_{\mathrm{PA}_{1}},U_{1}),\cdots,f(X_{\mathrm{PA}_{K}},U_{K}))$, where each $X_{k}:=f(X_{\mathrm{PA}_{k}},U_{k})$ for $k\in[K]$ (Pearl,, 2009).
Recent advances in deep learning-based causal inference (Chao et al.,, 2023; Shimizu,, 2023), particularly a class of methods integrating structural causal models (Pearl,, 2009) with generative models, have demonstrated the effectiveness of diffusion models for answering causal queries. However, spatial heterogeneity and temporal dependencies in unmeasured confounders undermine the validity of existing DCMs and BDCMs for genuine causal inference. For example, in environmental studies, pollution levels may vary significantly across regions (spatial heterogeneity) and exhibit serial correlation over time (temporal dependence), which standard causal models often fail to capture adequately. To overcome this fundamental limitation, we propose a Spatio-Temporal Dynamic Structural Causal Model (ST-DSCM) based on the Backdoor Criterion. We commence by establishing several essential definitions. Due to the complexity of the model, some symbols are mixed up in the article, but the overall readability remains unaffected.
2.1 Partially Functional Spatio-Temporal Dynamic Structural Causal Model
Suppose there is a spatio-temporal dataset $\boldsymbol{X}:=\{X_{k}\}_{k\in[K]}$ containing $K$ variables across $n$ regions over $J$ time points. Using $i$ to index regions and $j$ for time points, $X_{k}=(X_{ijk})_{n\times J}$ represent the value of the $k$-th variable $X_{k}$ over the time points and regions. We use a DAG to characterize the causal relationships among $\{X_{k}\}_{k\in[K]}$.
Consider a DAG $\mathcal{G}$ with nodes $[K]$ and a topologically sorted order such that each node $k$ has the $X_{k}$ as the random variable. Let $\mathcal{C}_{1}$ and $\mathcal{C}_{2}\subseteq[K]$ be two distinct sets of nodes with unobserved confounders, where $\mathcal{C}_{1}$ designates a set of unobserved explanatory nodes and $\mathcal{C}_{2}$ denotes a set of unobserved explained nodes. For $h=1,2$, let $\boldsymbol{X}_{\mathcal{C}_{h}}:=\{\boldsymbol{X}_{{\mathcal{C}_{h}}ij}\}_{i\in[n],j\in[J]}:=\{X_{qij}\}_{i\in[n],j\in[J],q\in\mathcal{C}_{h}}$, $\boldsymbol{U}_{\mathcal{C}_{h}}:=\{\boldsymbol{U}_{{\mathcal{C}_{h}}ij}\}_{i\in[n],j\in[J]}:=\{U_{qij}\}_{i\in[n],j\in[J],q\in\mathcal{C}_{h}}$.
$\boldsymbol{X}_{{\mathcal{C}_{2}}ij}=\boldsymbol{\Gamma}_{i}\boldsymbol{G}\left(\boldsymbol{X}_{{\mathcal{C}_{1}}ij}\right)+\boldsymbol{U}_{{\mathcal{C}_{2}}ij},$ (1)
where $\boldsymbol{\Gamma}_{i}$ is a $\dim(\boldsymbol{X}_{{\mathcal{C}_{2}}ij})\times\dim(\boldsymbol{X}_{{\mathcal{C}_{1}}ij})$ structural coefficient matrix, which are permitted to vary across regions; $\boldsymbol{G}(\cdot)=(g_{1}(.),\cdots,g_{q}(.))^{\top}$ is $q\times 1$ nonzero vector-valued function with differentiable functions $g_{1},\cdots,g_{q}$, and $q\geq\dim(\boldsymbol{X}_{{\mathcal{C}_{1}}ij})$. And to capture the temporal dependence, we let $\boldsymbol{X}_{{\mathcal{C}_{1}}i}=(\boldsymbol{X}_{{\mathcal{C}_{1}}i1}^{\top},\cdots,\boldsymbol{X}_{{\mathcal{C}_{1}}iJ}^{\top})^{\top}$ and $\boldsymbol{U}_{{\mathcal{C}_{2}}i}=(\boldsymbol{U}_{{\mathcal{C}_{2}}i1}^{\top},\cdots,\boldsymbol{U}_{{\mathcal{C}_{2}}iJ}^{\top})^{\top}$, then, the covariance matrices $\mathrm{Var}(\boldsymbol{X}_{{\mathcal{C}_{1}}i})$ and $\mathrm{Var}(\boldsymbol{U}_{{\mathcal{C}_{2}}i})$ of $\boldsymbol{X}_{{\mathcal{C}_{1}}i}$ and $\boldsymbol{U}_{{\mathcal{C}_{2}}i}$ can be expressed as
$\mathrm{Var}(\boldsymbol{X}_{{\mathcal{C}_{1}}i})=\boldsymbol{D}_{{\mathcal{C}_{1}}i}\otimes\mathrm{Var}(\boldsymbol{X}_{{\mathcal{C}_{1}}}),\ \mathrm{Var}(\boldsymbol{U}_{{\mathcal{C}_{2}}i})=\boldsymbol{D}_{{\mathcal{C}_{2}}i}\otimes\mathrm{Var}(\boldsymbol{U}_{\mathcal{C}_{2}}),$ (2)
where $\boldsymbol{D}_{{\mathcal{C}_{h}}i}(h\in\{1,2\})$ are the $J\times J$ adjacent time covariance matrices, $\mathrm{Var}(\cdot)$ represent between-variable covariances. In order to establish a rigorous framework for the temporal adjacency structure, the conditional autoregressive (CAR) model (Besag,, 1974) is adopted with the forms
$\boldsymbol{D}_{{{\mathcal{C}_{h}}i}}=\left(\boldsymbol{I}_{J}-\rho_{{{\mathcal{C}_{h}}i}}\boldsymbol{H}_{{{\mathcal{C}_{h}}i}}\right)^{-1},$ (3)
where $\rho_{{{\mathcal{C}_{h}}i}}(h\in\{1,2\})$ are adjacent time association parameters, and $\boldsymbol{H}_{{{\mathcal{C}_{h}}i}}(h\in\{1,2\})$ are $J\times J$ adjacency matrices in which the element $h_{jl}=1$ implies that time $l$ is adjacent to time $j$ and otherwise $h_{jl}=0$.
However, real-world applications often present significant challenges in elucidating causal relationships among variables observed under heterogeneity. To address these complexities, we extend the ST-DSCM by incorporating functional random variables $\boldsymbol{X}_{k}(t)$, leading to a Partially Functional Spatio-Temporal Dynamic Structural Causal Model (PFST-DSCM)  [^2] For example, Figure 2 is a PFST-DSCM with 33 exogenous and endogenous nodes, where nodes $X_{28}$, $X_{29}$ and $X_{30}$ are unmeasured confounders with spatial heterogeneity and temporal dependencies, $Y_{1}(t),Y_{2}(t),Y_{3}(t)$ are functional nodes, and $X_{4},\cdots,X_{21}$ are the corresponding base expansion nodes.. We adopt a basis expansion framework to achieve dimensionality reduction in the functional space via a set of orthogonal basis functions $\{\mathbf{b}_{1},\cdots,\mathbf{b}_{K_{n}}\}\in\mathbb{R}^{T\times K_{n}}$. Let
$X_{ijm}=\int\mathbf{b}_{m}(t)X_{ij}(t)\ \mathrm{d}t,$ (4)
for $i\in[n],j\in[J],m\in[K_{n}]$, which are used as nodes within the ST-DSCM, and detailed mathematical derivations are provided in Appendix A.
Notably, our model operates in settings where both observational data and causal structures are available, enabling it to answer observational, interventional, and counterfactual queries. All parameters in the structural model are assumed to be known in this context; in practice, they can be estimated using methods from Song et al., (2012) and Tang et al., (2017).
2.2 Partially Functional Dynamic Backdoor Diffusion-based Causal Model
We now present the PFD-BDCM, a model designed to handle causal queries under the PFST-DSCM framework, which explicitly accounts for unmeasured confounders with spatial heterogeneity and temporal dependence. The model employs an encoder-decoder architecture to enable causal reasoning across multi-resolution variables.
Let $\mathcal{B}_{k}$ be the set of backdoor nodes [^3] A set of node $\mathcal{B}$ satisfies backdoor criterion (Pearl et al.,, 2016) for tuple $(X,Y)$ in DAG $\mathcal{G}$ if no node in $\mathcal{B}$ is a descendant of $X$ and $\mathcal{B}$ blocks all paths between $X$ (cause) and $Y$ (outcome) which contains an arrow into $X$. of $k$, and $\boldsymbol{X}_{\mathcal{B}_{k}}:=\{X_{l}\}_{l\in\mathcal{B}_{k}}$ represent the variables on $\mathcal{B}_{k}$. Here, the set of parent nodes $\mathrm{PA}_{k}$ satisfies the definition of backdoor nodes, that is, $\mathrm{PA}_{k}\subseteq\mathcal{B}_{k}$, we will use $\mathcal{B}_{k}$ instead of $\mathrm{PA}_{k}$. We assume that the unobserved random variables are jointly independent (Markovian SCM), and the Partially Functional Spatio-Temporal Dynamic Directed Acyclic Graph (PFST-DDAG) $\mathcal{G}$ is the graph induced by PFST-DSCM $\mathcal{M}$. Every PFST-DSCM $\mathcal{M}$ entails a unique joint observational distribution satisfying the causal Markov assumption: $p(\mathbf{X})=\prod_{k=1}^{K}p(X_{k}|X_{\mathcal{B}_{k}})$.
The PFD-BDCM’s data-generating process is formalized as: $\{\mathbf{X}_{k}:X_{ijk}=f_{ij}(\mathbf{X}_{\mathcal{B}_{k}},\mathbf{U}_{k})\}_{k\in[K]}$. The encoder $g$ maps $(\mathbf{X}_{k},\mathbf{X}_{\mathcal{B}_{k}})$ to a latent variable $\mathbf{Z}_{k}:=g(\mathbf{X}_{k},\mathbf{X}_{\mathcal{B}_{k}})$, which captures information of the unmeasured confounders $\mathbf{U}_{k}$. The decoder $h$ reconstructs $\mathbf{X}_{k}$ as $\hat{\mathbf{X}}_{k}=h(\mathbf{Z}_{k},\mathbf{X}_{\mathcal{B}_{k}})$. Perfect reconstruction $\hat{\mathbf{X}}_{k}=\mathbf{X}_{k}$ implies that $h$ approximates the true structural function $f$. We next detail the architecture and training of PFD-BDCM, and then describe its use in causal query answering.